# Anti-bosonic current in thermal spin pumping

### v1

@@ Abstract
$=\int_{0}^{\infty}\frac{\omega_{R}^{n}\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}+\int_{-\infty}^{0}\frac{\omega_{L}^{n}\mathrm{DOS}(|\omega_{L}|)d|\omega_{L}|}{e^{\beta|\omega_{L}|}-1}$
@@
$\frac{1}{e^{\beta|\omega_{L}|}-1}=\frac{1}{e^{-\beta\omega_{L}}-1}.$ (22)
@@
$\mathcal{L}_{0}=$ $2G\left|S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}\right|\left(\int_{0}^{\infty}\frac{\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}+\int_{-\infty}^{0}\frac{\mathrm{DOS}(\omega)d\omega}{e^{-\beta\omega}-1}\right)$
$\mathcal{L}_{1}=$ $2G\left(S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}\right)\left(\int_{0}^{\infty}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}+\int_{-\infty}^{0}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{-\beta\omega}-1}\right)$
but subtractive in $\mathcal{L}_{1}$, analogous to the transport of electrons and holes in semiconductors (see Table 1). While band structure of semiconductor is weakly depends on temperature for as long as $k_{B}T\ll\epsilon_{F}$ is a lot smaller than the Fermi energy $\epsilon_{F}$, the magnon spectrum of ferrimagnetic Gd3Fe5O12 is strongly depends on temperature, as illustrated in Fig. 3)
@@
Boson Fermion
$R$-magnon ($\omega_{R}$ branch) Electron (conduction band)
@@
$L$-magnon ($\omega_{L}$ branch) Hole (valence band)
@@
High temperature spin pumping. We first consider temperatures higher than the compensation temperature $T>T_{B}$. In this regime, the net magnetization $S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}$ is positive. Thus $\omega_{R}=\omega_{1}$ and $\omega_{L}=\omega_{2}$. R-magnon dominate spin conductance at high temperatures because the number of L-magnon remains significantly smaller.
$\left(\frac{n_{L}}{n_{R}}\right)_{T>T_{B}}=\left(\frac{\mathcal{L}_{0L}}{\mathcal{L}_{0R}}\right)_{T>T_{B}}=\frac{\int_{\omega_{\mathrm{min}}}^{-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\frac{\mathrm{DOS}\left(\omega_{L}\right)d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{0}^{\infty}\frac{\mathrm{DOS}\left(\omega_{R}\right)d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
@@
In this temperature regime, the L-magnon contribution is still much smaller
$\left(\frac{\mathcal{L}_{1L}}{\mathcal{L}_{1R}}\right)_{T>T_{B}}=\frac{\int_{\omega_{\mathrm{min}}}^{-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\frac{\omega_{L}\mathrm{DOS}\left(\omega_{L}\right)d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{0}^{\infty}\frac{\omega_{R}\mathrm{DOS}\left(\omega_{R}\right)d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
$\simeq\frac{-8\left|\frac{S_{z}^{\mathrm{Fe}}}{S_{z}^{\mathrm{Gd}}}\right|^{\frac{3}{2}}}{3\zeta\left(\frac{5}{2}\right)}\int^{-\beta\lambda S^{\mathrm{Gd}}_{z}}_{0}\frac{\left[\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+x\right]x^{\frac{1}{2}}dx}{2\pi^{2}\left[e^{\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+x}-1\right]}\ll 1.$ (24)
Thus, R-magnon dominate $\mathcal{L}_{1}$ and the spin Seebeck coefficient $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ at high temperatures, as depicted in Fig. 4. We can see that R and L-magnons contributions are substractive in $\mathcal{L}_{1}$.
Figure 4: (a) The interface spin injection conductance $\mathcal{L}_{0}$.R-magnon dominates transport at high temperatures, while L-magnon becomes dominant at low temperatures. (b) The thermal spin current $\mathcal{L}_{1}$ exhibits an additional compensation temperature $T_{A}$, where L-magnon contribution exactly cancels R-magnon contribution. (c) The ratio $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ determine the measured spin Seebeck voltage ($V_{\mathrm{SSE}}$). Our theoretical result is in agreement with experiment data from Ref. Cramer et al., 2017.
Low temperature spin pumping. We now consider temperatures lower than the compensation temperature $T<T_{B}$, where $\omega_{R}=\omega_{2}$ and $\omega_{L}=\omega_{1}$. Thus, the magnon gap shifts to $\omega_{R}$. Thus, reducing number of R-magnon
$\left(\frac{n_{L}}{n_{R}}\right)_{T<T_{B}}=\left(\frac{\mathcal{L}_{0L}}{\mathcal{L}_{0R}}\right)_{T<T_{B}}=\frac{\int_{\omega_{\mathrm{min}}}^{0}\frac{\mathrm{DOS}(\omega_{L})d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}^{\infty}\frac{\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
$=\frac{\frac{2}{\sqrt{\pi}}\left|\frac{S_{z}^{\mathrm{Gd}}}{S_{z}^{\mathrm{Fe}}}\right|^{\frac{3}{2}}}{\mathrm{Li}_{\frac{3}{2}}\left(e^{\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\right)}\int^{\beta\lambda S^{\mathrm{Fe}}_{z}}_{0}\frac{x^{\frac{1}{2}}dx}{e^{x}-1}$
@@
$\left(\frac{\mathcal{L}_{1L}}{\mathcal{L}_{1R}}\right)_{T<T_{B}}=\frac{\int_{\omega_{\mathrm{min}}}^{0}\frac{\omega_{L}\mathrm{DOS}(\omega_{L})d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{-\lambda\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|}^{\infty}\frac{\omega_{R}\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
@@
Here $x_{0}=-\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)$. A compensation temperature $T_{A}$ occurs when the total $\mathcal{L}_{1R}$ and $\mathcal{L}_{1L}$ cancel each others, as illustrated in Fig.4b. The value of $\mathcal{L}_{1}$ becomes negative for $T_{A}<T<T_{B}$.Numerical evaluation of the integrals yields $T_{A}/T_{B}=0.3$, which aligns well with experimental values reported in Ref. Geprägs et al., 2016 ($T_{A}=80$ and $T_{B}=256.5$) and Ref. Cramer et al., 2017 ($T_{A}=72$ and $T_{B}\simeq 250$).
Conclusion. In summary, we explored the mechanisms behind thermal spin pumping in GdIG, emphasizing the competition of right-handed and left-handed magnons. As magnon gap appears at low temperature, R-magnon transport is suppressed and replaced by L-magnon. The theoretical framework developed highlights the crucial roles of antibosonic spin-wave dispersion in determining thermal spin pumping.
By deriving the temperature-dependent spin wave spectrum of GdIG and its influence on spin and heat response constants ($\mathcal{L}_{0}$ and $\mathcal{L}_{1}$), we demonstrated how antiferromagnetic magnons dynamics influence interface spin conductance and spin Seebeck coefficient. Our analysis underscores the significance of temperature-dependent magnons contributions, highlighting their competitive effects in thermal spin pumping. We identified distinct transport regimes where R-magnon dominate at high temperatures, while L-magnon become relevant at low temperatures. This dual contribution leads to notable features such as compensation points where thermal spin currents vanish. The compensation temperatures thermal spin pumping of GdIG were quantitatively captured.
The developed theoretical approach offers a comprehensive framework to understand thermal spin pumping phenomena in complex magnetic materials. These results can serve as a foundation for future experimental and computational studies aimed at optimizing spintronic devices based on thermal spin transport.

### v2

@@ 3 Results and Discussion
$=\int_{0}^{\infty}\frac{\omega_{R}^{n}\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}+\int_{-\infty}^{0}\frac{\omega_{L}^{n}\mathrm{DOS}(|\boldsymbol{\omega}_{L}|)d|\boldsymbol{\omega}_{L}|}{e^{\beta|\boldsymbol{\omega}_{L}|}-1}$
@@
$\frac{1}{e^{\beta|\boldsymbol{\omega}_{L}|}-1}=\frac{1}{e^{-\beta\omega_{L}}-1}.$ (15)
@@
$\frac{\mathcal{L}_{0}}{2G\left|S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}\right|}=\int_{0}^{\infty}\frac{\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}+\int_{-\infty}^{0}\frac{\mathrm{DOS}(\omega)d\omega}{e^{-\beta\omega}-1}$
but subtractive in $\mathcal{L}_{1}$
$\frac{\mathcal{L}_{1}}{2G\left(S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}\right)}=\int_{0}^{\infty}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{\beta\omega}-1}+\int_{-\infty}^{0}\frac{\omega\mathrm{DOS}(\omega)d\omega}{e^{-\beta\omega}-1}$
analogous to the transport of electrons and holes in semiconductors (see Table 1). While band structure of semiconductor is weakly depends on temperature for as long as $k_{B}T\ll\epsilon_{F}$ is a lot smaller than the Fermi energy $\epsilon_{F}$, the magnon spectrum of ferrimagnetic Gd3Fe5O12 is strongly depends on temperature, as illustrated in Fig. 3)
@@
Bosonic transport Fermionic transport
$R$-magnon Electron
($\omega_{R}$ branch) (conduction band)
@@
$L$-magnon Hole
($\omega_{L}$ branch) (valence band)
@@
3.1 High temperature spin pumping
We first consider temperatures higher than the compensation temperature $T>T_{B}$. In this regime, the net magnetization $S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}$ is positive. Thus $\omega_{R}=\omega_{1}$ and $\omega_{L}=\omega_{2}$. For small $k$, the magnon branches are
$\omega_{R}$ $\simeq$ $\displaystyle\frac{S^{\mathrm{Fe}}_{z}D_{\mathrm{YIG}}k^{2}}{S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}}$
$\omega_{L}$ $\simeq$ $-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+\displaystyle\frac{S^{\mathrm{Gd}}_{z}D_{\mathrm{YIG}}k^{2}}{S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}}.$ (16)
R-magnon dominate spin conductance at high temperatures because the number of L-magnon remains significantly smaller.
$\left(\frac{n_{L}}{n_{R}}\right)_{T>T_{B}}=\left(\frac{\mathcal{L}_{0L}}{\mathcal{L}_{0R}}\right)_{T>T_{B}}=\frac{\int_{-\lambda S^{\mathrm{Fe}}_{z}}^{-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\frac{\mathrm{DOS}\left(\omega_{L}\right)d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{0}^{\infty}\frac{\mathrm{DOS}\left(\omega_{R}\right)d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
@@ 3.1 High temperature spin pumping
In this temperature regime there is $|\boldsymbol{\omega}|$ spectrum crossing. However, the thermal spin pumping remains finite because the L-magnon contribution is still much smaller
$\left(\frac{\mathcal{L}_{1L}}{\mathcal{L}_{1R}}\right)_{T>T_{B}}=\frac{\int_{-\lambda S^{\mathrm{Fe}}_{z}}^{-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\frac{\omega_{L}\mathrm{DOS}\left(\omega_{L}\right)d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{0}^{\infty}\frac{\omega_{R}\mathrm{DOS}\left(\omega_{R}\right)d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
$\simeq\frac{-8\left|\frac{S_{z}^{\mathrm{Fe}}}{S_{z}^{\mathrm{Gd}}}\right|^{\frac{3}{2}}}{3\zeta\left(\frac{5}{2}\right)}\int^{-\beta\lambda S^{\mathrm{Gd}}_{z}}_{0}\frac{\left[\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+x\right]x^{\frac{1}{2}}dx}{2\pi^{2}\left[e^{\beta\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+x}-1\right]}\ll 1.$
Thus, R-magnon dominate $\mathcal{L}_{1}$ and the spin Seebeck coefficient $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ at high temperatures, as depicted in Fig. 4.
Figure 4: (a) The interface spin injection conductance $\mathcal{L}_{0}$.R-magnon dominates transport at high temperatures, while L-magnon becomes dominant at low temperatures. (b) The thermal spin current $\mathcal{L}_{1}$ exhibits an additional compensation temperature $T_{A}$, where L-magnon contribution exactly cancels R-magnon contribution. (c) The ratio $\mathcal{L}_{1}/(\mathcal{L}_{0}T)$ determine the measured spin Seebeck voltage ($V_{\mathrm{SSE}}$). Our theoretical result is in agreement with experiment data from Ref. [25].
3.2 Low temperature spin pumping
We now consider temperatures lower than the compensation temperature $T<T_{B}$, where $\omega_{R}=\omega_{2}$ and $\omega_{L}=\omega_{1}$. For small $k$, the magnon branches are
$\omega_{R}$ $\simeq$ $\lambda\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|-\frac{S^{\mathrm{Gd}}_{z}D_{\mathrm{YIG}}k^{2}}{\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|},$
$\omega_{L}$ $\simeq$ $-\frac{S^{\mathrm{Fe}}_{z}D_{\mathrm{YIG}}k^{2}}{\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|}.$ (19)
The magnon gap shifts to $\omega_{R}$. Thus, reducing number of R-magnon
$\left(\frac{n_{L}}{n_{R}}\right)_{T<T_{B}}$ $=$ $\left(\frac{\mathcal{L}_{0L}}{\mathcal{L}_{0R}}\right)_{T<T_{B}}=\frac{\int_{-\lambda S^{\mathrm{Fe}}_{z}}^{0}\frac{\mathrm{DOS}(\omega_{L})d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{\lambda\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|}^{\infty}\frac{\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}}$ (20)
$=$ $\frac{\frac{2}{\sqrt{\pi}}\left|\frac{S_{z}^{\mathrm{Gd}}}{S_{z}^{\mathrm{Fe}}}\right|^{\frac{3}{2}}}{\mathrm{Li}_{\frac{3}{2}}\left(e^{-\beta\lambda\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)}\right|}\int^{\beta\lambda S^{\mathrm{Fe}}_{z}}_{0}\frac{x^{\frac{1}{2}}dx}{e^{x}-1}$
@@ 3.2 Low temperature spin pumping
$\left(\frac{\mathcal{L}_{1L}}{\mathcal{L}_{1R}}\right)_{T<T_{B}}=\frac{\int_{-\lambda S^{\mathrm{Fe}}_{z}}^{0}\frac{\omega_{L}\mathrm{DOS}(\omega_{L})d\omega_{L}}{e^{-\beta\omega_{L}}-1}}{\int_{-\lambda\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|}^{\infty}\frac{\omega_{R}\mathrm{DOS}(\omega_{R})d\omega_{R}}{e^{\beta\omega_{R}}-1}}$
@@
Here $x_{0}=\beta\lambda\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|$. The asymptote value for low temperature is
$\lim_{T\to 0}\mathcal{L}_{1}=\frac{3G\zeta\left(\frac{5}{2}\right)\left(k_{B}T\right)^{\frac{5}{2}}\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|^{\frac{5}{2}}}{4\pi^{\frac{3}{2}}D_{0}^{\frac{3}{2}}\left(S^{\mathrm{Fe}}_{z}\right)^{\frac{3}{2}}}.$ (23)
@@
Although there is no $|\boldsymbol{\omega}|$ crossing when $T<T_{B}$, a compensation temperature $T_{A}$ occurs when the total $\mathcal{L}_{1R}$ and $\mathcal{L}_{1L}$ cancel each others, as illustrated in Fig.4b. At $T_{A}$, the thermal spin current carried by right-handed and left-handed magnons cancel each other. Reciprocally, the net heat current carried by both magnon modes in under finite spin accumulation is also zero due to Onsager relation in matriks Eq. 1. The value of $\mathcal{L}_{1}$ becomes negative for $T_{A}<T<T_{B}$. Numerical evaluation of the integrals yields $T_{A}/T_{B}=0.3$, which aligns well with experimental values reported in Ref. [26] ($T_{A}=80$ K and $T_{B}=256.5$ K) and Ref. [25] ($T_{A}=72$ K and $T_{B}\simeq 250$ K).
Our findings contrast with the interpretation presented in Ref. [10], where the compensation temperature is attributed to a spectral crossing of $|\boldsymbol{\omega}_{R}|$ and $|\boldsymbol{\omega}_{L}|$ obtained from a full multi-sublattice exchange model. In that framework, the vanishing of the spin current is interpreted as a direct consequence of spectral symmetry. In our model, however, we show that such a spectral crossing is neither a necessary nor sufficient condition for compensation. Instead, we identify the compensation temperature $T_{A}$ as arising from a precise cancellation between the thermally weighted contributions of right- and left-handed magnons, governed by their distinct frequency branches and temperature-dependent populations. Our results therefore offer an alternative interpretation of the origin of compensation and highlight the role of magnon statistics rather than spectral degeneracy alone. This distinction provides theoretical clarity and suggests that compensation may occur under broader conditions than previously assumed.
4 Conclusion
In this work, we investigated the mechanisms of thermal spin pumping in gadolinium iron garnet (GdIG) by analyzing the distinct contributions of left-handed and right-handed antiferromagnetic magnons. Using a two-magnetic-lattice model with simplified Fe–Fe and Fe–Gd exchange interactions, we derived the temperature-dependent spin-wave spectrum and evaluated the corresponding spin transport coefficients $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$.
We treated right-handed and left-handed magnons as positive and negative frequency branches, respectively. This approach is conceptually analogous to electron and hole transport in semiconductors, where carriers in opposite bands contribute with opposite signs to current. At high temperatures, right-handed magnons dominate the thermal spin current due to their higher population. As the temperature decreases, their contribution becomes suppressed due to the opening of a magnon gap, while left-handed magnons emerge as the primary carriers.
This competition leads to a compensation temperature, at which the thermally driven spin current vanishes. Notably, we found that this compensation point does not coincide with the crossing of $|\boldsymbol{\omega}_{R}|$ and $|\boldsymbol{\omega}_{L}|$ spectra. Furthermore, we observed that spin pumping remains finite at temperatures where such crossings occur. This behavior highlights the fact that compensation results from the integrated statistical contributions of each mode, rather than from spectral symmetry alone.
Our findings provide an alternative interpretation to that of Ref. [10], which attributes compensation to a spectral crossing in a detailed multi-sublattice exchange model. In contrast, our simplified two-sublattice framework captures key features of spin pumping behavior in a more analytically transparent form. This model offers both qualitative and quantitative insights into mode-resolved spin transport in ferrimagnets.
The theoretical framework developed in this work contributes a physically intuitive and computationally efficient tool for analyzing thermal spin pumping. It is well suited for guiding experimental interpretation and for informing the design of spintronic devices based on chiral magnon dynamics and thermally driven spin currents.
@@ 4 Conclusion
Appendix A Gd-Gd exchange contribution to compensation temperatures
Here we estimate the contribution of Gd-Gd exchange, characterized by $I_{\mathrm{Gd}}\ll\lambda\ll I$, to the compensation temperatures using the following Hamiltonian that includes magnetic field
$\mathcal{H}=$ $-\mu_{0}\textbf{H}\cdot 2\mu_{B}\left(\textbf{S}^{\mathrm{Fe}}+\textbf{S}^{\mathrm{Gd}}\right)-\frac{I}{2}\sum_{j}\textbf{S}^{\mathrm{Fe}}_{j}\cdot\textbf{S}^{\mathrm{Fe}}_{j+1}$
$+\lambda\sum_{j}\textbf{S}^{\mathrm{Gd}}_{j}\cdot\textbf{S}^{\mathrm{Fe}}_{n}-\frac{I_{\mathrm{Gd}}}{2}\sum_{j}\textbf{S}^{\mathrm{Gd}}_{j}\cdot\textbf{S}^{\mathrm{Gd}}_{j+1},$
where $\mu_{B}$ is the Bohr magneton. The statistical average is as follows
$S^{\mathrm{Gd}}_{z}$ $=$ $3S^{\mathrm{Gd}}B_{S^{\mathrm{Gd}}}\left(S^{\mathrm{Gd}}\frac{2\mu_{B}\mu_{0}H+I_{\mathrm{Gd}}S^{\mathrm{Gd}}_{z}-\lambda S^{\mathrm{Fe}}_{z}}{k_{B}T}\right)$
$S^{\mathrm{Fe}}_{z}$ $=$ $S^{\mathrm{Fe}}B_{S^{\mathrm{Fe}}}\left(S^{\mathrm{Fe}}\frac{2\mu_{B}\mu_{0}H+IS^{\mathrm{Fe}}_{z}-\lambda S^{\mathrm{Gd}}_{z}}{k_{B}T}\right).$
A.1 Contribution on $T_{C}$
For high temperatures, $S^{\mathrm{Gd}}_{z}$ and $S^{\mathrm{Fe}}_{z}$ can be linearised as follows
$\left[\begin{array}[]{cc}1-\frac{I^{\mathrm{Gd}}S^{\mathrm{Gd}}\left(S^{\mathrm{Gd}}+1\right)}{k_{B}T}&\frac{\lambda}{k_{B}T}\\ \frac{\lambda}{k_{B}T}&1-\frac{IS^{\mathrm{Fe}}\left(S^{\mathrm{Fe}}+1\right)}{3k_{B}T}\end{array}\right]\left[\begin{array}[]{c}S^{\mathrm{Gd}}_{z}\\ S^{\mathrm{Fe}}_{z}\end{array}\right]$
$\approx\left[\begin{array}[]{c}\frac{2\mu_{B}\mu_{0}HS^{\mathrm{Gd}}\left(S^{\mathrm{Gd}}+1\right)}{k_{B}T}\\ \frac{2\mu_{B}\mu_{0}HS^{\mathrm{Fe}}\left(S^{\mathrm{Fe}}+1\right)}{3k_{B}T}\end{array}\right]$
The Curie temperature occurred when the determinant of the matrix is zero.
$T_{C}$$\approx$ $\frac{IS^{\mathrm{Fe}}\left(S^{\mathrm{Fe}}+1\right)}{3k_{B}}+\frac{3\lambda^{2}}{k_{B}IS^{\mathrm{Fe}}\left(S^{\mathrm{Fe}}+1\right)}\left(1+\frac{3I_{\mathrm{Gd}}}{I}\right)$
A.2 Contribution on $T_{B}$
$S^{\mathrm{Gd}}_{z}$ on zero magnetic field can be linearized as follows
$S^{\mathrm{Gd}}_{z}(T)\approx S^{\mathrm{Gd}}\left(S^{\mathrm{Gd}}+1\right)\frac{I_{\mathrm{Gd}}S^{\mathrm{Gd}}_{z}-\lambda S^{\mathrm{Fe}}_{z}}{k_{B}T}$
$\approx-\lambda S^{\mathrm{Fe}}_{z}\frac{S^{\mathrm{Gd}}\left(S^{\mathrm{Gd}}+1\right)}{k_{B}T}\left(1+\frac{I_{\mathrm{Gd}}S^{\mathrm{Gd}}\left(S^{\mathrm{Gd}}+1\right)}{k_{B}T}\right).$
The modified compensation temperature $T_{B}$ as follows
00 0 $=$ $S^{\mathrm{Gd}}_{z}(T_{B})+S^{\mathrm{Fe}}_{z}$
00 0 $=$ $1-\lambda\frac{S^{\mathrm{Gd}}\left(S^{\mathrm{Gd}}+1\right)}{k_{B}T_{B}}\left(1+\frac{I_{\mathrm{Gd}}S^{\mathrm{Gd}}\left(S^{\mathrm{Gd}}+1\right)}{k_{B}T}\right)$
such that
$T_{B}\approx\frac{\lambda S^{\mathrm{Gd}}\left(S^{\mathrm{Gd}}+1\right)}{k_{B}}\left(1+\frac{I_{\mathrm{Gd}}}{\lambda}\right)$
A.3 Contribution on $T_{A}$
The Landau–Lifshitz equations of motion for the spins become
$\frac{d\textbf{S}^{\mathrm{Fe}}_{j}}{dt}$ $=$ $\textbf{S}^{\mathrm{Fe}}_{j}\times\left(\frac{I}{2}\left(\textbf{S}^{\mathrm{Fe}}_{j+1}+\textbf{S}^{\mathrm{Fe}}_{j-1}\right)-\lambda\textbf{S}^{\mathrm{Gd}}_{j}\right)$
$\frac{d\textbf{S}^{\mathrm{Gd}}_{j}}{dt}$ $=$ $\textbf{S}^{\mathrm{Gd}}_{j}\times\left(-\lambda\textbf{S}^{\mathrm{Fe}}_{j}+\frac{I_{\mathrm{Gd}}}{2}\left(\textbf{S}^{\mathrm{Gd}}_{j+1}+\textbf{S}^{\mathrm{Gd}}_{j-1}\right)\right).$
The spin-wave spectrum for small $k$ is
$i\frac{d}{dt}\left[\begin{array}[]{c}S^{\mathrm{Fe}}_{+}\\ S^{\mathrm{Gd}}_{+}\end{array}\right]=W\left[\begin{array}[]{c}S^{\mathrm{Fe}}_{+}\\ S^{\mathrm{Gd}}_{+}\end{array}\right],$
$W=\left[\begin{array}[]{cc}IS^{\mathrm{Fe}}_{z}\frac{(ka)^{2}}{2}-\lambda S^{\mathrm{Gd}}_{z}&\lambda S^{\mathrm{Fe}}_{z}\\ \lambda S^{\mathrm{Gd}}_{z}&I_{\mathrm{Gd}}S^{\mathrm{Gd}}_{z}\frac{(ka)^{2}}{2}-\lambda S^{\mathrm{Fe}}_{z}\end{array}\right]$
the dispersion relation are
$\frac{k^{2}a^{2}}{2}=\frac{\omega\left(\omega+\lambda\left(S_{z}^{\mathrm{Fe}}+S_{z}^{\mathrm{Gd}}\right)\right)}{\left(IS_{z}^{\mathrm{Fe}}+I_{\mathrm{Gd}}S_{z}^{\mathrm{Gd}}\right)+\lambda\left(I\left(S_{z}^{\mathrm{Fe}}\right)^{2}+I_{\mathrm{Gd}}\left(S_{z}^{\mathrm{Gd}}\right)^{2}\right)}$
The minimum frequency is now
$\omega_{\mathrm{min}}\simeq-\lambda S^{\mathrm{Fe}}_{z}\left(1+\frac{I_{\mathrm{Gd}}S_{z}^{\mathrm{Gd}}}{IS_{z}^{\mathrm{Fe}}}\left(\frac{S_{z}^{\mathrm{Gd}}}{S_{z}^{\mathrm{Fe}}}-1\right)\right).$
$I_{\mathrm{Gd}}$ modifies the spin wave dispersion, which in turn determine $T_{A}$. For $T>T_{B}$,
$\omega_{R}$$\approx$ $\displaystyle\frac{S^{\mathrm{Fe}}_{z}D_{\mathrm{YIG}}k^{2}}{S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}}\left(1+\frac{I_{\mathrm{Gd}}}{I}\left(\frac{S^{\mathrm{Gd}}_{z}}{S^{\mathrm{Fe}}_{z}}\right)^{2}\right)$
$\omega_{L}$ $\simeq$ $-\lambda\left(S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right)+\displaystyle\frac{S^{\mathrm{Gd}}_{z}D_{\mathrm{YIG}}k^{2}}{S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}}\left(1+\frac{I_{\mathrm{Gd}}}{I}\right).$
Meanwhile, for $T<T_{B}$,
$\omega_{R}$$\approx$ $\lambda\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|-\frac{S^{\mathrm{Gd}}_{z}D_{\mathrm{YIG}}k^{2}}{\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|}\left(1+\frac{I_{\mathrm{Gd}}}{I}\right),$
$\omega_{L}$$\approx$ $-\frac{S^{\mathrm{Fe}}_{z}D_{\mathrm{YIG}}k^{2}}{\left|S^{\mathrm{Fe}}_{z}+S^{\mathrm{Gd}}_{z}\right|}\left(1+\frac{I_{\mathrm{Gd}}}{I}\left(\frac{S^{\mathrm{Gd}}_{z}}{S^{\mathrm{Fe}}_{z}}\right)^{2}\right).$